AM3 -Esercitazione 8 A.A. 2003-2004

Prof. Luigi Chierchia, Dott. Laura Di Gregorio 30 aprile 2004

Esercizio 1

Calcolare l’integrale

∫∫_{S} xy dxdy

esteso al dominio S che è limitato dall’asse OX e dal semicerchio superiore (x − 2)^{2} + y^{2} = 1.

Esercizio 2

Calcolare

∫∫_{D} \frac{x^{2}}{y} e^{xy} dx dy

dove

D = \{(x, y) ∈ ℝ^{2} : \frac{1}{2x} ≤ y ≤ \frac{1}{x}, 2x^{2} ≤ y ≤ 3x^{2}\}.